# Aligning Text, Images, and 3D Structure Token-by-Token

### v1

@@ A.3 Tokenization
Text. We employ an off-the-shelf text tokenizer from Llama-3.2 grattafiori2024llama3herdmodels with a vocabulary size of $128,\!000$ for tokenization.
@@ A.4 Task sequence formation
After tokenizing all three modalities, we construct sequences tailored to the specific task, which are then used to train Kyvo. In this section, we present examples of complete sequences for each of the all tasks across the four datasets.
@@
[…]
[BOS][SCENE-START][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.80 -2.00 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]brown[MATERIAL]metal[SHAPE]cylinder[LOCATION]-1.35 2.65 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]purple[MATERIAL]rubber[SHAPE]sphere[LOCATION]-0.90 -2.20 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]rubber[SHAPE]cylinder[LOCATION]-2.70 -2.90 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.25 -1.15 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]red[MATERIAL]metal[SHAPE]cube[LOCATION]2.15 -2.65 0.35[OBJECT-END][SCENE-END][OUTPUT-SEP][IMAGE-START]<image-tokens>[IMAGE-END][EOS]
@@
[BOS][IMAGE-START]<image-tokens>[IMAGE-END][OUTPUT-SEP][SCENE-START][OBJECT-START][SIZE]small[COLOR]brown[MATERIAL]rubber[SHAPE]sphere[LOCATION]-2.50 0.30 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]blue[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.95 1.60 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]-0.85 0.60 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][SCENE-END][EOS]
[…]
[BOS][IMAGE-START]<image-tokens>[IMAGE-END][SCENE-START][OBJECT-START][SIZE]small[COLOR]brown[MATERIAL]rubber[SHAPE]sphere[LOCATION]-2.50 0.30 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]blue[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.95 1.60 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]-0.85 0.60 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][SCENE-END][TEXT-START]Change the brown object to have purple color[TEXT-END]
[OUTPUT-SEP][IMAGE-START]<image-tokens>[IMAGE-END][SCENE-START][OBJECT-START][SIZE]small[COLOR]purple[MATERIAL]rubber[SHAPE]sphere[LOCATION]-2.50 0.30 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]blue[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.95 1.60 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]-0.85 0.60 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][SCENE-END][EOS]
[…]
[BOS][IMAGE-START]<image-tokens>[IMAGE-END][SCENE-START][OBJECT-START][SIZE]small[COLOR]brown[MATERIAL]rubber[SHAPE]sphere[LOCATION]-2.50 0.30 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]blue[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.95 1.60 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]-0.85 0.60 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][SCENE-END][TEXT-START]What size is the rubber sphere?[TEXT-END][OUTPUT-SEP][TEXT-START]small[TEXT-END][EOS]
@@
[BOS][SCENE-START][OBJECT-START][SHAPE]<$v_{1}^{1}$,$v_{2}^{1}$,...,$v_{512}^{1}$>[LOCATION]2.10 0.15 -0.75[POSE]0.00 0.00 -2.10[OBJECT-END][OBJECT-START][SHAPE]<$v_{1}^{2}$,$v_{2}^{2}$,...,$v_{512}^{2}$>[LOCATION]-1.25 2.85 0.05[POSE]0.00 0.00 1.85[OBJECT-END][OBJECT-START][SHAPE]<$v_{1}^{3}$,$v_{2}^{3}$,...,$v_{512}^{3}$>[LOCATION]0.45 -2.35 -1.50[POSE]0.00 0.00 0.60[OBJECT-END][SCENE-END][OUTPUT-SEP][IMAGE-START]<image-tokens>[IMAGE-END][EOS]
@@
[BOS][IMAGE-START]<image-tokens>[IMAGE-END][OUTPUT-SEP][SCENE-START][OBJECT-START][SHAPE]<$v_{1}^{1}$,$v_{2}^{1}$,...,$v_{512}^{1}$>[LOCATION]1.15 -2.85 0.40[POSE]0.00 0.00 -1.75[OBJECT-END][OBJECT-START][SHAPE]<$v_{1}^{2}$,$v_{2}^{2}$,...,$v_{512}^{2}$>[LOCATION]-0.35 2.10 -0.50[POSE]0.00 0.00 2.45[OBJECT-END][OBJECT-START][SHAPE]<$v_{1}^{3}$,$v_{2}^{3}$,...,$v_{512}^{3}$>[LOCATION]2.75 0.15 -1.60[POSE]0.00 0.00 -0.25[OBJECT-END][SCENE-END][EOS]
@@ Appendix B Additional qualitative examples
Rendering. We present qualitative results for the rendering task on CLEVR and ObjaWorld with complex shapes in Figures 9 and  10, respectively. Results on ObjaWorld with explicit shape representations are shown in Figure 11. The model takes only the structured 3D scene representation as input and predicts the corresponding image tokens. These tokens are then decoded using the VQGAN decoder, which reconstructs the image by mapping them from token-space to pixel-space. For each example, we also provide the ground truth image rendered using Blender, allowing direct comparison with the model-generated output. As observed, the model effectively captures the 3D scene structure based solely on the JSON input and accurately synthesizes the corresponding image. Notably, Figure 11 demonstrates the model’s ability to integrate multiple information sources: it successfully maps spatial arrangements from JSON scene descriptions and visual properties from asset sequences to generate coherent, realistic pixel-space representations.
@@
In Figure 13 we show examples of recognition task on ObjaWorld with explicit shape representations on unseen scenes. Specifically, the model takes a single images as input and reconstructs the full 3D geometry per object and infers each object’s 3D position and pose to accurately reconstruct the 3D scene. As can be seen from Figure 13, Kyvo accurately recovers object geometries and spatial layouts via our structured, object-centric 3D modality. In contrast, Trellis xiang2024structured often merges two objects into one (e.g. the second example) or hallucinates shapes (e.g., blue can in the first example) and misaligned layouts.
@@ Appendix C Evaluation
In this section we provide more details on the evaluation strategies that we adopt. We discuss the evaluation method for the three modalities.
@@
  • ObjaWorld: Match shape with pose constraint (predicted pose within ±$\pm 0.15$ radians)
@@
[…]
@@
[…]
@@
[…]
@@
26:  $\texttt{JaccardIndexAvg}\leftarrow\texttt{JaccardIndex}/|\text{PS}|$
@@ C.2 Images
Score: The score takes a binary value of “00” or “$1$” and signifies the complete correctness of an image generation. The user is asked to provide a score of “$1$” only if the user believes that all the objects in the scene were accurately generated and accurately placed spatially. If the generated image has any differences with the groundtruth, e.g. a cube was not generated correctly, the user provides a score of “00". The score value is independent of any other model involved in the comparison and solely depends on the model under consideration and the groundtruth.
[…]
@@
C.3 Text
@@
True/False $,08$ 080 False
@@ D.1 Image VQGAN architecture
In this section, we provide additional implementation details for the VQGAN architecture employed in our experiments to represent images. Figure 18 illustrates the complete network architecture, with output shapes shown for each layer given an input of shape ($1$, $3$, $256$, $256$). Our VQGAN configuration uses a $1024$-entry codebook with $256$-dimensional embeddings, trained on $256\times 256$ resolution images. The encoder produces quantized embeddings of shape ($1$, $256$, $16$, $16$), yielding $16\times 16=256$ discrete tokens per image that correspond to learned codebook entries. For each dataset, we initialize the VQGAN model with ImageNet pretrained weights from esser2021taming , then fine-tune on the respective training set for 100 epochs. This fine-tuned model serves as both the encoder for converting images to token sequences and the decoder for reconstructing images from predicted tokens during evaluation.
@@
Figure 18: VQGAN Architecture. We show the detailed architecture of the VQGAN model that we use to train a domain-specific codebook. We show the output shapes for an input shape ($1$, $3$, $256$, $256$).
D.2 Encoding of numbers

### v2

@@ A.3 Tokenization
Text. We employ an off-the-shelf text tokenizer from Llama-3.2 [grattafiori2024llama3herdmodels] with a vocabulary size of 128,000128,\!000 for tokenization.
@@ A.4 Task sequence formation
After tokenizing all three modalities, we construct sequences tailored to the specific task, which are then used to train Kyvo. In this section, we present examples of complete sequences for each of the tasks across the four datasets.
@@
Rendering: $\text{3D}\rightarrow\text{Image}$
[BOS][SCENE-START][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.80 -2.00 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]brown[MATERIAL]metal[SHAPE]cylinder[LOCATION]-1.35 2.65 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]purple[MATERIAL]rubber[SHAPE]sphere[LOCATION]-0.90 -2.20 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]rubber[SHAPE]cylinder[LOCATION]-2.70 -2.90 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.25 -1.15 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]red[MATERIAL]metal[SHAPE]cube[LOCATION]2.15 -2.65 0.35[OBJECT-END][SCENE-END][OUTPUT-SEP]
[IMAGE-START]<image-tokens>[IMAGE-END][EOS]
@@
[BOS][IMAGE-START]<image-tokens>[IMAGE-END][OUTPUT-SEP][SCENE-START]
[OBJECT-START][SIZE]small[COLOR]brown[MATERIAL]rubber[SHAPE]sphere[LOCATION]-2.50 0.30 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]blue[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.95 1.60 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]-0.85 0.60 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][SCENE-END][EOS]
Instruction-Following: $(\text{Image},\text{3D},\text{Text}_{\text{I}})\rightarrow(\text{Image},\text{3D})$
[BOS][IMAGE-START]<image-tokens>[IMAGE-END][SCENE-START][OBJECT-START][SIZE]
small[COLOR]brown[MATERIAL]rubber[SHAPE]sphere[LOCATION]-2.50 0.30 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]blue[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.95 1.60 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]-0.85 0.60 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][SCENE-END][TEXT-START]Change the brown object to have purple color[TEXT-END]
[OUTPUT-SEP][IMAGE-START]<image-tokens>[IMAGE-END][SCENE-START][OBJECT-START][SIZE]
small[COLOR]purple[MATERIAL]rubber[SHAPE]sphere[LOCATION]-2.50 0.30 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]blue[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.95 1.60 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]-0.85 0.60 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][SCENE-END][EOS]
Question-Answering: $(\text{Image},\text{3D},\text{Text}_{\text{Q}})\rightarrow\text{Text}_{\text{A}}$
[BOS][IMAGE-START]<image-tokens>[IMAGE-END][SCENE-START][OBJECT-START][SIZE]
small[COLOR]brown[MATERIAL]rubber[SHAPE]sphere[LOCATION]-2.50 0.30 0.35[OBJECT-END][OBJECT-START][SIZE]large[COLOR]blue[MATERIAL]metal[SHAPE]cylinder[LOCATION]2.95 1.60 0.70[OBJECT-END][OBJECT-START][SIZE]large[COLOR]red[MATERIAL]metal[SHAPE]cylinder[LOCATION]-0.85 0.60 0.70[OBJECT-END][OBJECT-START][SIZE]small[COLOR]green[MATERIAL]metal[SHAPE]sphere[LOCATION]0.85 1.85 0.35[OBJECT-END][SCENE-END][TEXT-START]What size is the rubber sphere?[TEXT-END][OUTPUT-SEP][TEXT-START]small[TEXT-END][EOS]
@@
[IMAGE-START]<image-tokens>[IMAGE-END][EOS]
@@
[BOS][IMAGE-START]<image-tokens>[IMAGE-END][OUTPUT-SEP][SCENE-START]
[…]
@@ Appendix B Additional qualitative examples
3D Tokenization. We present additional qualitative results for our 3D tokenization scheme. Fig.˜11 provides additional examples of how the 3D tokens are effective for reconstructions, as shown in Figure 4 (b) in the paper. Fig.˜12 provides additional examples of how the 3D tokens are effective for decoding, as shown in Figure 4 (c) in the paper.
Figure 11: Reconstruction examples for Trellis-based 3D VQ-VAE. VQ-VAE reconstructions, shown from multiple views, compared to ground truth for unseen Objaverse assets.
Figure 12: Decoding examples for Trellis-based 3D VQ-VAE on unseen Objaverse assets. (a) Reconstruction of 3D shape encoding using Llama 3.2 as decoder. (b) Rendering from 3D shape encoding using Llama 3.2 as decoder.
Rendering. We present qualitative results for the rendering task on CLEVR and ObjaWorld with complex shapes in Figures 13 and  14, respectively. Results on ObjaWorld with explicit shape representations are shown in Figure 15. The model takes only the structured 3D scene representation as input and predicts the corresponding image tokens. These tokens are then decoded using the VQGAN decoder, which reconstructs the image by mapping them from token-space to pixel-space. For each example, we also provide the ground truth image rendered using Blender, allowing direct comparison with the model-generated output. As observed, the model effectively captures the 3D scene structure based solely on the JSON input and accurately synthesizes the corresponding image. Notably, Figure 15 demonstrates the model’s ability to integrate multiple information sources: it successfully maps spatial arrangements from JSON scene descriptions and visual properties from asset sequences to generate coherent, realistic pixel-space representations.
@@
In Figure 17 we show examples of recognition task on ObjaWorld with explicit shape representations on unseen scenes. Specifically, the model takes a single images as input and reconstructs the full 3D geometry per object and infers each object’s 3D position and pose to accurately reconstruct the 3D scene. As can be seen from Figure 17, Kyvo accurately recovers object geometries and spatial layouts via our structured, object-centric 3D modality. In contrast, Trellis [xiang2024structured] often merges two objects into one (e.g. the second example) or hallucinates shapes (e.g., blue can in the first example) and misaligned layouts.
@@ Appendix C Evaluation
In this section we provide more details on the evaluation strategies that we adopt. We discuss the evaluation method for the three primary modalities as well as for the 3D shape encodings.
@@
  • ObjaWorld: Match shape with pose constraint (predicted pose within $\pm 0.15$ radians)
@@
7:  matchedFlags $\leftarrow$ Boolean array of length $|\texttt{GObjs}|$, initialized to False
@@
11:   for $j=1$ to $|\texttt{GObjs}|$ do
12:    if $\neg\texttt{matchedFlags}[j]$ and $\texttt{attributesMatch}(\texttt{p},\texttt{GObjs}[j])$ and $\texttt{dist}(\texttt{p.coords},\texttt{GObjs}[j].\texttt{coords})<\tau$ then
@@
23:  $\texttt{FN}\leftarrow\texttt{FN}+\text{count}(\texttt{matchedFlags}=\texttt{False})$
@@
[…]
@@ C.2 Images
Score: The score takes a binary value of “0” or “11” and signifies the complete correctness of an image generation. The user is asked to provide a score of “11” only if the user believes that all the objects in the scene were accurately generated and accurately placed spatially. If the generated image has any differences with the groundtruth, e.g. a cube was not generated correctly, the user provides a score of “0". The score value is independent of any other model involved in the comparison and solely depends on the model under consideration and the groundtruth.
Rank: The rank takes a value from $\{$11”, “22”, …, “N}\}. The user is expected to rank the N images from “11" to “N" by comparing the generation quality among them. If the user is unable to decide between two images, then we allow equal rank assignment to more than one image, e.g. if two models equally perform for a given scene, they get the same rank value.
@@
C.3 3D assets
For the same reasons as above, we use human evaluations to assess the quality of 3D assets when it comes to the 3D tokenizer. In Table 1 in the main paper, we use human evaluations to quantify the effect of auxiliary reconstruction losses on the reconstruction quality of the 3D VQ-VAE, and report mean rank. In Table 2 in the main paper, we use human evaluations to compare the reconstruction quality of our Trellis-based 3D VQ-VAE with that of SAR3D, and report mean rank. In both cases, the human evaluation is facilitated using an anonymized, shuffled interface as described above, and only relative rank is assessed and reported.
C.4 Text
@@
True/False $8,080$ False
@@ D.1 Image VQGAN architecture
In this section, we provide additional implementation details for the VQGAN architecture employed in our experiments to represent images. Figure 22 illustrates the complete network architecture, with output shapes shown for each layer given an input of shape ($1$, 33, $256$, $256$). Our VQGAN configuration uses a $1024$-entry codebook with $256$-dimensional embeddings, trained on $256\times 256$ resolution images. The encoder produces quantized embeddings of shape ($1$, $256$, $16$, $16$), yielding $16\times 16=256$ discrete tokens per image that correspond to learned codebook entries. For each dataset, we initialize the VQGAN model with ImageNet pretrained weights from [esser2021taming], then fine-tune on the respective training set for 100 epochs. This fine-tuned model serves as both the encoder for converting images to token sequences and the decoder for reconstructing images from predicted tokens during evaluation.
@@
Figure 22: VQGAN Architecture. We show the detailed architecture of the VQGAN model that we use to train a domain-specific codebook. We show the output shapes for an input shape ($1$, 33, $256$, $256$).
D.2 3D VQ-VAE training
In Section 3.2.1 of the main paper we describe how we train a 3D VQ-VAE to compress the $64^{3}\times 8$ slats to a dense $8^{3}\times 128$ latent. Here, we provide some additional details on the training. A slat contains $\sim\!20$k sparse voxels (in a $64^{3}$ grid, with vectors of dimension $8$ at active positions). We train the 3D VQ-VAE with these slats in 3 steps: (i) it densifies the sparse tensor onto a $64^{3}\!\times 8$ grid, (ii) encodes that grid with a 3-D convolutional U-Net down to an $8^{3}\!\times 128$ latent volume, and (iii) vector-quantizes every latent with an $8192$-entry codebook.
Architecture.
The original Trellis SLAT encoder and decoder remain frozen and provide supervision. Sparse latents $(z_{i},p_{i})$ are first rasterised into a zero-initialised $64^{3}\!\times 8$ grid; using a learnable padding vector instead of zeros showed no measurable gain and is therefore omitted. The dense grid is processed by a 3D U-Net that downsamples as $64^{3}\!\to 32^{3}\!\to 16^{3}\!\to 8^{3}$ with channel widths $(32,128,512,1024)$. Each $8^{3}$ cell outputs a $128$-dim vector, quantized to the nearest of $8192$ code vectors updated by exponential moving average ($\tau{=}0.99$). Gradients flow to the encoder through the straight-through estimator.
Optimization.
We train for 200200k steps on $\sim 168$k Sketchfab assets from Objaverse (further details provided in Sec.˜A.1) with a batch size of $8$. We use an AdamW optimizer ($\beta_{1}{=}0.9,\beta_{2}{=}0.999$) with a constant learning-rate of $3{\times}10^{-4}$, no weight decay, mixed precision, and adaptive gradient clipping.
The training objective combines four terms:
$\mathcal{L}={}$ $\underbrace{\|x-\hat{x}\|_{2}^{2}}_{\text{dense-SLAT recon}}+\beta\,\underbrace{\|\,z-\mathrm{sg}[e]\|_{2}^{2}}_{\text{commit}}$ (1)
$\quad+\lambda_{\text{KL}}\,D_{\text{KL}}+\gamma\,\mathcal{L}_{\text{render}}$
where $x\in\mathbb{R}^{64^{3}\times 8}$ is the rasterised SLAT voxel, $\hat{x}$ its VQ–VAE reconstruction, $\mathrm{sg}[\cdot]$ denotes the stop-gradient operator. In the commitment term, $z\in\mathbb{R}^{8^{3}\times 128}$ denotes the pre-quantization output of the encoder, while $e$ is the corresponding vector-quantized output selected from the codebook. The $\mathcal{L}_{\text{render}}$ follows TRELLIS:
[…]
computed between images rendered from ground-truth images $I$ and rendering from Gaussian reconstructions $\hat{I}$. This is the auxiliary pixel-space reconstruction loss discussed in the main paper. We set $\beta=0.25,\;\lambda_{\text{KL}}=10^{-6},\;\gamma=0.1$.
The resulting codebook is used to represent objects using $8^{3}=512$ discrete tokens, enabling unified processing with structured 3D representations for autoregressive modeling. We employ bi-directional attention within shape token sequences for full intra-sequence connectivity when training the LLM.
Codebook usage. Figure 23 shows codebook utilization across 5000 training assets (log-scale, decreasing order). The heavy-tailed distribution indicates effective learning: frequent codes capture common 3D primitives while rare codes encode geometric variations. Active utilization of the whole codebook indicates it is not over-parameterised. Additionally, increasing codebook size did not help empirically indicating it is not under-parameterised.
Figure 23: Codebook usage. Histogram of usage counts (log scale) for the $8192$ latent codes on $5000$ assets used for training. The heavy-tailed distribution reveals a compact core of frequently reused codes that capture ubiquitous 3D primitives, while the long tail represents rarer geometric variations. In total, all of the vocabulary is exercised at least once, indicating that the codebook is not over-parameterised. Additionally, increasing codebook size did not help empirically indicating it is not under-parameterised.
The full training script and code for 3D VQ-VAE are included in the code.zip file. The code was built on top of Trellis [xiang2024structured] code.
D.3 Encoding of numbers